Johns father is 61 years old. The difference between his age 4 years ago and johns age in 1 year is equal to Johns age now. How old is John?
Answer by ankor@dixie-net.com(22740)

let j = John's present age
:
Johns father is 61 years old.
The difference between his age 4 years ago and johns age in 1 year is equal to Johns age now.
61 - 4 - (j + 1) = j
57 - j - 1 = j
57 - 1 = j + j
56 = 2j
j = 56/2
j = 28 yrs